what would be the domain of this equation:
The square root of 9-4x. All is included under the radical!
thanks

Found 2 solutions by Edwin McCravy, venugopalramana:
Answer by Edwin McCravy(20086)

what would be the domain of this equation: 
The square root of 9-4x. All is included under the 
radical! thanks

           ______
          Ö9 - 4x

An even root, such as square root, 4th root, 6th root,
etc. (BUT NOT an odd root, such as cube root, 5th 
root, 7th root, etc) gives a real result when and only 
when its radicand (what's under the radical) is not 
negative, that is, greater than or equal to 0.

This is an even root, so for the domain we must require
that its radicand be greater than or equal to 0:

         9 - 4x ³ 0

            -4x ³ -9

              x £ 9/4

Graph of domain on a number line:


< =========================·--------
                          9/4

Domain in set builder notation {x|x £ 9/4}

Domain in interval notation:  (-¥, 9/4] 


--------------------------------------------

Note: if it were an odd root instead of an 
even root the domain would always be

(-¥, ¥)

Edwin

Answer by venugopalramana(3286)

Y = SQRT(9-4X)
THE FUNCTION IS NOT DEFINED IF 9-4X UNDER RADICAL SIGN IS NEGATIVE.
HENCE 9-4X>=0
9>=4X
9/4=2.25>=X
X<=2.25
SO DOMAIN OF THE FUNCTION IS ALL REAL VALUES >=2.25...THAT IS
(-INFINITY,2.25]